# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ Abstract
Text-to-image diffusion models (DMs) have achieved remarkable success in image generation. However, concerns about data privacy and intellectual property remain due to their potential to inadvertently memorize and replicate training data. Recent mitigation efforts have focused on identifying and pruning weights responsible for triggering replication, based on the assumption that memorization can be localized. Our research assesses the robustness of these pruning-based approaches. We demonstrate that even after pruning, minor adjustments to text embeddings of input prompts are sufficient to re-trigger data replication, highlighting the fragility of these defenses. Furthermore, we challenge the fundamental assumption of memorization locality, by showing that replication can be triggered from diverse locations within the text embedding space, and follows different paths in the model. Our findings indicate that existing mitigation strategies are insufficient and underscore the need for methods that truly remove memorized content, rather than attempting to suppress its retrieval. As a first step in this direction, we introduce a novel adversarial fine-tuning method that iteratively searches for replication triggers and updates the model to increase robustness. Through our research, we provide fresh insights into the nature of memorization in text-to-image DMs and a foundation for building more trustworthy and compliant generative AI.
@@ 1 Introduction
Generating high-quality images with diffusion models (DMs) enjoys great popularity. However, undesired memorization of training data in text-to-image DMs [37, 2] poses significant risks to privacy and intellectual property, as it can favor the unintended replication of sensitive or copyrighted content during inference. In response, various detection and mitigation strategies have been proposed [37, 43, 45, 32]. Most existing mitigation techniques either aim to identify and filter out highly memorized samples during training [37, 32] or modify inputs at inference time [37, 45, 32] to reduce memorization-induced data replication. While the training-based methods require computationally expensive retraining, the inference-time methods are limited to models behind APIs, as users of open-source models can easily disable these mechanisms by altering the source code.
@@
To overcome both limitations, recent approaches [14, 3] observe that the text prompts of memorized images elicit distinct activation patterns in the DMs. Based on these activations, the methods prune a small set of weights, effectively reducing the risk of verbatim data replication, while preserving overall image quality. However, since these methods work with a single prompt per memorized image, it remains an open question whether they prevent the replication of memorized images through other inputs. We search for Diffusion Memorization (Dori [Uncaptioned image]) beyond the prompt space by crafting adversarial embeddings—text embeddings different from the memorized prompts—that trigger memorized image generations. Adversarial embeddings allow us to recover supposedly removed memorized data after pruning (see Fig. 1, left), revealing that pruning merely conceals memorization.
Initial effectiveness of pruning-based methods for mitigating memorization is attributed to a locality phenomenon, a property of the model to store memorized data in a small (local) set of memorization weights. Intuitively, if this property holds, then pruning memorization weights would prevent generation of memorized data. In this work, we challenge the locality phenomenon and question whether locality is real or a misinterpretation of the early success of pruning-based methods. We investigate the input space and the model’s weights and find little support for it. Memorization seems to be spread out, as there exist multiple adversarial embeddings that cause replication of the same data point, with the DM following different paths during generation, see Fig. 1 (right). Similarly, the activation patterns and memorization weights identified for the same memorized image vary across different inputs that trigger its replication, further undermining the notion of locality.
Accordingly, a robust memorization removal method should avoid the pitfall of assuming locality. To this end, we build on adversarial embeddings and develop a novel adversarial fine-tuning approach to completely erase memorized samples from text-to-image DMs. Our approach is inspired by adversarial training [40, 10, 26], which iteratively generates adversarial examples to train robust models. While prior methods have focused on parameter pruning or input adjustments, our approach directly modifies the model’s parameters to eliminate memorization. In contrast to pruning-based mitigation techniques, our method achieves reliable removal and remains robust against adversarial embeddings designed to circumvent mitigation.
@@
    We reveal that existing weight-pruning methods merely conceal memorization in text-to-image DMs rather than truly erase memorized content from a model.
@@
    We challenge the assumption that memorization is local, demonstrating that it fails to hold across both the input space and the model parameters.
@@
    As a more robust defense, we propose a novel adversarial fine-tuning scheme that permanently mitigates memorization in already trained DMs.
Figure 1: Left: 1 Without mitigation, the DM closely replicates the training sample. 2 Mitigation strategies, such as pruning memorization neurons with NeMo [14] or Wanda [3], prevent replication for the memorized prompt, thereby suggesting successful removal. Yet, 3 adversarial embeddings Refer to caption still trigger replication. Right: While pruning alters the generation trajectory for the original memorized prompt (blue), adversarial embeddings steer denoising along alternative paths (red) that still lead to the memorized content, unaffected by the pruning-based mitigation.
@@ 2 Background and Related Work
In this section, we explore the core principles of text-to-image generation using DMs and examine research focused on the critical issue of unintended memorization within this context.
@@ 2.1 Text-to-Image Generation with Diffusion Models
Diffusion models [38, 15] (DMs) are a class of generative models trained by gradually corrupting training images by adding Gaussian noise and training a model $\bm{\epsilon}_{\bm{\theta}}$ to predict the noise that has been added. Once trained, DMs generate new images by starting from pure noise ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and progressively denoising it. At each time step $t=T,\ldots,1$, the model $\bm{\epsilon}_{\bm{\theta}}$ predicts the noise $\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t,{\bm{y}})$ needed for the denoising step. In the domain of text-to-image generation, the denoising process is guided by a text prompt ${\bm{p}}$, which is transformed into a text embedding ${\bm{y}}$ by a text encoder.
During training, a time step $t\sim\mathcal{U}(1,T)$ and a noise vector $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ are randomly sampled to create a noisy image ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ based on the training image ${\bm{x}}_{0}$. The amount of noise added is controlled by a noise scheduler $\bar{\alpha}_{t}$, for which there are multiple choices [38, 27, 20, 19]. The training objective for the noise predictor $\bm{\epsilon}_{\bm{\theta}}$ is then to predict the noise $\bm{\epsilon}$ that has been added:
Training and generating samples with DMs can be computationally expensive. The latent DM framework [33] reduces this burden by operating in a lower-dimensional latent space instead of the pixel space. This latent space is learned by a separately trained variational autoencoder [22, 41] that encodes images into compact representations and decodes generated latents back to the image space.
@@ 2.2 Memorization in Diffusion Models
Definition. In the context of generative models, memorization [9, 8] can manifest as the model reproducing portions of its training data, such as closely replicating a particular sample. Specifically, verbatim memorization (VM) describes cases when a training image is reliably generated by the model with almost a pixel-perfect match. Template memorization (TM) is a more relaxed notion, in which we require that only parts of the image are closely replicated, such as the background of an image or a specific object [43].
@@
Memorization in DMs. Recent work has demonstrated that DMs—especially text-to-image models [33, 35]—are also prone to such unintended memorization [37, 2, 18, 12, 6, 25, 7, 46, 17], raising concerns around privacy and intellectual property. Since then, multiple methods have been developed to detect data replication [43, 45, 32, 23]. While many of these techniques rely on the availability of training prompts to identify memorized content, another line of research detects memorization even in the absence of training prompts, focusing instead on identifying specific memorized images [25, 16].
@@
Mitigation. Memorization in DMs can either be prevented during training or by intervening in the generation process at inference time. Existing training-time mitigation techniques either adjust the training data by removing duplicates [2, 37] or reject training samples for which the model indicates signs of memorization [45, 32, 5]. However, since re-training large DMs is expensive, inference-time mitigation strategies are crucial for already trained models. These mitigation strategies adjust the input tokens [37], update the text embeddings [45], change the cross-attention scores [32], or guide the noise prediction away from memorized content [4]. However, these methods offer no permanent mitigation, increase the inference time, and can easily be turned off for locally deployed models.
@@
Local Pruning-Based Mitigation. More permanent solutions have focused on identifying and removing the weights responsible for triggering data replication. Hintersdorf et al. [14] developed NeMo, a localization algorithm to detect memorization neurons within the cross-attention value layers of DMs, of which all weights are pruned. More specifically, NeMo first conducts an out-of-distribution detection to identify neurons with high absolute activations under memorized prompts and reduces the set of identified neurons by checking their influence on data replication individually. Similarly, Chavhan et al. [3] applied Wanda [39]—a pruning technique originally developed for large language models—to locate and prune individual weights in the output fully-connected layers of cross-attention modules responsible for memorization. Wanda identifies weights by their weight importance, computed as the product between the weights and the activation norm. The method then prunes the top kof weights with the highest importance scores compared to scores computed on a null string. While both methods successfully avoid data replication triggered by memorized prompts, it remains an open question whether the memorized content is successfully erased from the model.
@@ 3 Breaking Pruning-Based Mitigation Methods
This section introduces our method for finding triggers of memorized content in text embeddings, which we use to critically assess the effectiveness of the two pruning-based mitigation methods NeMo and Wanda. Our analysis demonstrates that while both methods successfully eliminate data replication when faced with memorized prompts, they are highly vulnerable to adversarial embeddings.
@@ 3.1 Finding Dori [Uncaptioned image] With Adversarial Text Embeddings
[…]
@@ 3.2 Experimental Setup
Models and Datasets: We focus our investigation on Stable Diffusion v1.4 [33] and a set of $500$ memorized prompts [45, 43] from the LAION-5B [36] training dataset, in line with previous research [45, 32, 14, 3] on memorization in text-to-image DMs. More recent DMs are trained on more carefully curated and deduplicated datasets, which reduces the amount of memorization, as discussed in previous work [37, 44]. While results for TM prompts are included in Appx. E, the main paper focuses on VM prompts, addressing this arguably more critical form of memorization.
@@
Metrics: Following prior work [45, 14], we employ SSCD [30], a feature extractor commonly employed to detect and quantify copying behavior in DMs. To measure similarity between two images, we compute the cosine similarity between their SSCD feature embeddings. Higher values indicate a higher degree of content replication. All metrics are computed as the median of the maximum scores across ten generated images per memorized prompt or adversarial embedding. We vary the seeds for image generation and adversarial embedding optimization to avoid overfitting.
@@
To evaluate replication, we define $\text{SSCD}_{\text{Orig}}$ as the cosine similarity between generated images and their associated training image. Values above $0.7$ indicate that the memorized image is successfully generated [45]. Additionally, we use the similarity between images generated before and after mitigation techniques are applied, denoted by $\text{SSCD}_{\text{Gen}}$, to assess the effects of mitigation. We expect lower $\text{SSCD}_{\text{Gen}}$ scores when mitigation is successful.
Typically, memorized images are consistently replicated, regardless of the choice of the initial noise $\bm{\epsilon}$. Conversely, for any other input, images generated by a DM are diverse under varying initial noise. We quantify diversity as the average pairwise cosine similarity $\text{D}_{\text{SSCD}}$ between SSCD embeddings of images generated from the same input but different initial noise. Images generated after mitigation should exhibit greater diversity, indicated by lower $\text{D}_{\text{SSCD}}$ values.
@@
To assess image quality, we measure prompt alignment using CLIP [31] similarity $A_{\text{CLIP}}$, comparing each generated image with its corresponding textual prompt. Higher $A_{\text{CLIP}}$ scores indicate a stronger semantic alignment with the input prompt. We also compute the Fréchet Inception Distance (FID) [13] and Kernel Inception Distance (KID, reported in the Appendix) [1]. Both are evaluated on 30k prompts of the COCO dataset [24]—a standard benchmark for image generation [29]. Lower scores indicate improved image quality.
Mitigation Methods: We base our research on pruning-based mitigation methods, specifically NeMo [14] and Wanda [3], which are currently the only existing mitigation methods of this category. For both methods, we use the default hyperparameters specified in the respective papers. For NeMo, we set the threshold used to identify memorization neurons to $\tau_{\text{mem}}=0.428$ (mean neuron activations plus one standard deviation). For Wanda, we use a weight sparsity of $1\%$. Following the original evaluation settings, we identify and evaluate mitigation sample-wise for NeMo, whereas Wanda identifies the set of weights to be pruned for all memorized images at once using the first 10 timesteps.
@@
Adversarial Embedding Optimization: We initialize the adversarial embedding optimization with the text embeddings of the prompts of the memorized images. We then optimize each embedding for $50$ steps with a learning rate of $0.1$ using Adam [21] and a batch size of $8$.

### v2

# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Not Local

@@ Abstract
Text-to-image diffusion models (DMs) have achieved remarkable success in image generation. However, concerns about data privacy and intellectual property remain due to their potential to inadvertently memorize and replicate training data. Recent mitigation efforts have focused on identifying and pruning weights responsible for triggering verbatim training data replication, based on the assumption that memorization can be localized. We challenge this assumption and demonstrate that, even after such pruning, small perturbations to the text embeddings of previously mitigated prompts can re-trigger data replication, revealing the fragility of such defenses. Our further analysis then provides multiple indications that memorization is indeed not inherently local: (1) replication triggers for memorized images are distributed throughout text embedding space; (2) embeddings yielding the same replicated image produce divergent model activations; and (3) different pruning methods identify inconsistent sets of memorization-related weights for the same image. Finally, we show that bypassing the locality assumption enables more robust mitigation through adversarial fine-tuning. These findings provide new insights into the nature of memorization in text-to-image DMs and inform the development of more reliable mitigations against DM memorization.Contact: antoni.kowalczuk@cispa.de, dominik.hintersdorf@dfki.de, lukas.struppek@dfki.de
@@ 1 Introduction
Generating high-quality images with diffusion models (DMs) enjoys great popularity. However, undesired memorization and verbatim replication of training data in text-to-image DMs (Somepalli et al., 2023; Carlini et al., 2023) poses significant risks to privacy and intellectual property, as it can favor the unintended replication of sensitive or copyrighted data points during inference. In response, various detection and mitigation strategies have been proposed (Somepalli et al., 2023; Webster, 2023; Wen et al., 2024; Ren et al., 2024). Most existing mitigation techniques either aim to identify and filter out highly memorized samples during training (Somepalli et al., 2023; Ren et al., 2024) or modify inputs at inference time (Somepalli et al., 2023; Wen et al., 2024; Ren et al., 2024) to reduce memorization-induced data replication. While the training-based methods require computationally expensive retraining, the inference-time methods are limited to models behind APIs, as users of open-source models can easily disable these mechanisms by altering the source code.
@@
To overcome both limitations, recent approaches (Hintersdorf et al., 2024; Chavhan et al., 2024) observe that the text prompts of memorized images elicit distinct activation patterns in the DMs. Based on these activations, the methods prune a small set of weights, effectively reducing the risk of verbatim data replication, while preserving overall image quality. However, since these methods work with a single prompt per memorized image, it remains an open question whether they prevent the replication of memorized images through different inputs. We search for Diffusion Memorization (Dori [Uncaptioned image]) beyond the prompt space by crafting adversarial embeddings, i.e., text embeddings different from the memorized prompts, that trigger generation of memorized images. Such adversarial embeddings allow us to recover supposedly removed memorized data after pruning (see Fig.˜1, left), revealing that pruning merely conceals verbatim memorization. Rather than being limited to a subset of individual weights, memorization appears to be distributed throughout the model. For a single memorized data point, multiple adversarial embeddings can trigger its replication, with the DM following different generation paths, see Fig.˜1 (right). Similarly, the different activation patterns and memorization weights identified for the same memorized image vary across different inputs that trigger its replication, further undermining the notion of locality.
Abandoning the locality assumption, we develop the first memorization removal method effective against adversarial embeddings. We employ adversarial training (Szegedy et al., 2014; Goodfellow et al., 2014; Madry et al., 2018), which iteratively searches for adversarial embeddings that trigger replication, and pair it with full fine-tuning rather than pruning a subset of weights to achieve reliable removal of memorized data points.
@@
    We reveal that existing weight-pruning methods merely conceal memorization in text-to-image DMs rather than truly erase memorized individual data points from a model.
@@
    We challenge the assumption that memorization is local, demonstrating that locality fails to hold across (1) the text embedding space, (2) a model’s activations, and (3) its trained weights.
@@
    Finally, we introduce fine-tuning with adversarial text embeddings as a strong memorization removal method, demonstrating that memorization can be permanently mitigated in already trained DMs without relying on locality, paving the way for more refined methods.
Figure 1: Left:
1
Without mitigation, the DM closely replicates the training sample.
2
Mitigation strategies, such as pruning memorization neurons with NeMo (Hintersdorf et al., 2024) or Wanda (Chavhan et al., 2024), prevent replication for the memorized prompt, thereby suggesting successful removal. Yet,
3
adversarial embeddings Refer to caption still trigger replication. Right: While pruning alters the generation trajectory for the original memorized prompt (blue), adversarial embeddings steer denoising along alternative paths (red) that still lead to the memorized content, unaffected by the pruning-based mitigation.
@@ 2 Background and Related Work
In this section, we present the core principles of text-to-image generation using DMs and introduce research focused on unintended memorization of individual training data points. We also discuss the fundamental differences between mitigating memorization and concept unlearning.
@@ 2.1 Text-to-Image Generation with Diffusion Models
Diffusion models (Song & Ermon, 2020; Ho et al., 2020) (DMs) are a class of generative models trained by gradually corrupting training images by adding Gaussian noise and training a model $\bm{\epsilon}_{\bm{\theta}}$ to predict the noise that has been added. Once trained, DMs generate new images by starting from pure noise ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and progressively denoising it. At each time step $t=T,\ldots,1$, the model $\bm{\epsilon}_{\bm{\theta}}$ predicts the noise $\bm{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t,{\bm{y}})$ needed for the denoising step. In the domain of text-to-image generation, the denoising process is guided by a text prompt $\bm{p}$}, which is transformed into a text embedding $\bm{y}$} by a text encoder. We discuss more technical details on their training in Section˜E.1.
@@ 2.2 Memorization in Diffusion Models
Definition. In the context of generative models, memorization (Feldman & Zhang, 2020; Feldman, 2020) can manifest as the model reproducing portions of its training data, such as closely replicating a particular individual training sample. Specifically, verbatim memorization (VM) describes cases when a training image is reliably generated by the model with almost a pixel-perfect match. Template memorization (TM) is a more relaxed notion, in which only parts of the image are closely replicated, such as the background of an image or a specific object (Webster, 2023). Especially, verbatim generation of individual training data points has a detrimental effect on the trustworthy deployment of DMs, as it can lead to privacy leaks and copyright violations if sensitive and copyrighted data is included in the models’ training set.
@@
Memorization in DMs. Recent work has demonstrated that DMs, especially text-to-image models (Rombach et al., 2022; Saharia et al., 2022), are prone to unintended data point memorization (Somepalli et al., 2023; Carlini et al., 2023; Kadkhodaie et al., 2024; Gu et al., 2023; Chen et al., 2024b; Ma et al., 2024; Dar et al., 2023; Zhang et al., 2024a), raising concerns around privacy and intellectual property. Since then, multiple methods have been developed to detect data replication (Webster, 2023; Wen et al., 2024; Ren et al., 2024; Kriplani et al., ). While many of these techniques rely on the availability of training prompts to identify memorized content, another line of research detects memorization even in the absence of training prompts, focusing instead on identifying specific memorized images (Ma et al., 2024; Jiang et al., 2025).
@@
Mitigation. Memorization in DMs can either be prevented during training or by intervening in the generation process at inference time. Existing training-time mitigation techniques either adjust the training data by removing duplicates (Carlini et al., 2023; Somepalli et al., 2023) or reject training samples for which the model indicates signs of memorization (Wen et al., 2024; Ren et al., 2024; Chen et al., 2025). However, since re-training large DMs is expensive, inference-time mitigation strategies are crucial for already trained models. These mitigation strategies adjust the input tokens (Somepalli et al., 2023), update the text embeddings (Wen et al., 2024), change the cross-attention scores (Ren et al., 2024), or guide the noise prediction away from memorized content (Chen et al., 2024a). However, all these methods offer no permanent mitigation, increase the inference time, and can easily be turned off for locally deployed models.
@@
Local Pruning-Based Mitigation. More permanent solutions have focused on identifying and removing the weights responsible for triggering data replication. Hintersdorf et al. (2024) developed NeMo, a localization algorithm to detect memorization neurons within the cross-attention value layers of DMs, of which all weights are pruned. More specifically, NeMo first conducts an out-of-distribution detection to identify neurons with high absolute activations under memorized prompts and reduces the set of identified neurons by checking their influence on data replication individually. Similarly, Chavhan et al. (2024) applied Wanda (Sun et al., 2024), a pruning technique originally developed for large language models, to locate and prune individual weights in the output fully-connected layers of transformer blocks responsible for memorization. Wanda identifies weights by their weight importance, computed as the product between the weights and the activation norm. The method then prunes the top k% of weights with the highest importance scores compared to scores computed on a null string. While both methods successfully avoid data replication triggered by memorized prompts, it remains open whether the memorized data points are successfully erased from the model.
Memorization Mitigation vs. Concept Unlearning in DMs. Apart from the localization-based memorization mitigation techniques, one of the few approaches that try to remove information from DMs’ weights are concept unlearning methods (Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2024b) that are used for content moderation. Although these methods bear some methodological similarity, they pursue fundamentally different objectives. Concept unlearning targets the suppression of broad, high-level concepts, such as nudity or specific objects (e.g., cars), across all generations. In contrast, memorization mitigation (this work) seeks to remove the model’s ability to reproduce specific, individual memorized training data points. For example, mitigating verbatim generation of a particular memorized image of a car to protect copyright prevents the model from generating that exact image, but does not affect its capacity to generate cars in general. Concept unlearning, on the other hand, eliminates the model’s ability to generate any car, which is undesirable when we want to mitigate memorization of a specific data point but leave the model unchanged otherwise. Therefore, mitigating memorization, i.e., (verbatim) training data replication, although deceptively similar to concept removal, needs a different approach. In the next sections, we validate this empirically by showing that concept removal methods are indeed not suitable for mitigating the memorization of individual data points, which calls for stronger, tailored tools.
@@ 3 Breaking Pruning-Based Mitigation Methods
In this section, we highlight that pruning-based memorization mitigations only conceal memorization but fail to truly remove memorized images from DMs. Specifically, we show that even after applying these mitigations, we can still trigger the generation of memorized images through carefully crafted adversarial text embedding. This reveals that, despite the apparent mitigation of memorization under standard textual prompts, the underlying memorized images persist in the model weights.
@@ 3.1 Finding Dori [Uncaptioned image] With Adversarial Text Embeddings
To demonstrate that pruning-based memorization mitigation strategies (NeMo and Wanda) fail to truly remove memorized images from DMs, we develop a novel approach for generating adversarial text embeddings that can still trigger the verbatim reproduction of supposedly removed memorized data points. Instead of relying on natural-language prompts, we use unconstrained continuous optimization in the text embedding space to uncover triggers capable of reconstructing memorized images, even after a mitigation has been applied. The existence of such adversarial triggers reveals that memorized content is still encoded in the model weights and can be verbatim extracted, which poses a significant privacy and copyright risk, especially for open-weight models.
Formally, let ${\bm{x}}_{\mathit{mem}}$ be a known memorized image and ${\bm{y}}_{\mathit{mem}}$ the text embedding for its prompt. After weight pruning using NeMo or Wanda, the model no longer replicates ${\bm{x}}_{\mathit{mem}}$ when conditioned on ${\bm{y}}_{\mathit{mem}}$, giving the impression that memorization has been successfully removed. To verify removal, we optimize an adversarial embedding ${\bm{y}}_{adv}$ initialized with ${\bm{y}}_{\mathit{mem}}$, using gradients of the standard diffusion loss $\mathcal{L}_{\mathit{DM}}$ (see Eq.˜3), with learning rate $\eta$ over multiple steps $i$ as:
${\bm{y}}_{\mathit{adv}}^{(i+1)}={\bm{y}}_{\mathit{adv}}^{(i)}-\eta\nabla_{{\bm{y}}_{\mathit{adv}}^{(i)}}\mathcal{L}_{\mathit{DM}}({\bm{x}}_{\mathit{mem}},\bm{\epsilon},{\bm{y}}_{\mathit{adv}}^{(i)},t,\bm{\theta}_{\mathit{NeMo/Wanda}})\text{,}$ (1)
where $\bm{\theta}_{\mathit{NeMo/Wanda}}$ are parameters of the DM after applying NeMo or Wanda to mitigate replication of ${\bm{x}}_{\mathit{mem}}$. Our goal is to find a final adversarial embedding ${\bm{y}}_{adv}$ that consistently triggers ${\bm{x}}_{\mathit{mem}}$, regardless of the initial noise, so we re-sample the noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ at each optimization step. Similarly, we re-sample the timesteps $t\sim\mathcal{U}(1,T)$ to ensure that ${\bm{y}}_{adv}$ reliably triggers ${\bm{x}}_{\mathit{mem}}$ during the iterative denoising generation process of the DM. A detailed formulation of the optimization procedure is provided in Alg.˜1 in Appx.˜F.
For comparison, we also evaluate UnlearnDiffAtk (Zhang et al., 2024c), a state-of-the-art attack designed to re-trigger forgotten concepts in the context of concept unlearning. As established above, concept unlearning fundamentally differs from the challenge of removing individual memorized training examples, yet UnlearnDiffAtk represents the closest available baseline and is therefore included for completeness. Our results (see Section˜G.1) show that UnlearnDiffAtk fails to re-generate memorized images following pruning-based mitigations. This highlights both the particular challenge of showing the limits of pruning-based memorization techniques and the necessity of our adversarial optimization strategy, which enables analysis beyond what was possible with prior methods.
@@ 3.2 Experimental Setup
Models and Datasets: We focus our investigation on Stable Diffusion v1.4 (Rombach et al., 2022) and a set of $500$ memorized prompts (Wen et al., 2024; Webster, 2023) from the LAION-5B (Schuhmann et al., 2022) training dataset, in line with previous research on memorization in text-to-image DMs (Wen et al., 2024; Ren et al., 2024; Hintersdorf et al., 2024; Chavhan et al., 2024). More recent DMs are trained on more carefully curated and deduplicated datasets, which reduces the amount of memorization, as discussed in previous work (Somepalli et al., 2023; Webster et al., 2023). Therefore, Stable Diffusion v1.4 is the only existing model for which a known set of memorized prompts is publicly available. As a result, it is currently not possible to conduct comparable memorization studies on other models. In the main paper, we focus on VM prompts as they represent the most critical form of memorization, but we additionally include results for TM prompts in Appx.˜H.
@@
Metrics: Following prior work (Wen et al., 2024; Hintersdorf et al., 2024), we employ SSCD (Pizzi et al., 2022), a feature extractor commonly employed to detect and quantify copying behavior in DMs. To measure similarity between two images, we compute the cosine similarity between their SSCD feature embeddings. Higher values indicate a higher degree of content replication. All metrics are computed as the median of the maximum scores across ten generated images per memorized prompt or adversarial embedding. We vary the seeds for image generation and adversarial embedding optimization to avoid overfitting.
@@
To evaluate replication, we define $\text{SSCD}_{\text{Orig}}$ as the cosine similarity between generated images and their associated training image. Values above $0.7$ indicate that the memorized image is successfully generated (Wen et al., 2024). Additionally, we use the similarity between images generated before and after mitigation techniques are applied, denoted by $\text{SSCD}_{\text{Gen}}$, to assess the effects of mitigation. We expect lower $\text{SSCD}_{\text{Gen}}$ scores when mitigation is successful.
While SSCD-based metrics capture the overall trend of replication, we also quantify the number of images that remain memorized even after mitigation. To this end, we define the Memorization Rate (MR) as the ratio of memorized images that the model can still replicate, i.e., those achieving an SSCD score above $0.7$ for at least one of the ten generations, sampled from different random seeds.
@@
To assess image quality, we measure prompt alignment using CLIP (Radford et al., 2021) similarity $A_{\text{CLIP}}$, comparing each generated image with its corresponding textual prompt. Higher $A_{\text{CLIP}}$ scores indicate a stronger semantic alignment with the input prompt. We also compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID (Binkowski et al., 2018), reported in the Appendix). Both these image quality metrics are evaluated on 30k prompts of the COCO dataset (Lin et al., 2014), a standard benchmark for image generation (Pavlov et al., 2023). Lower scores indicate improved image quality.
@@
Adversarial Embedding Optimization: We initialize the adversarial embeddings with the original text embeddings of the prompts for the memorized images. We then optimize each embedding for $50$ steps with a learning rate of $0.1$ using Adam (Kingma & Ba, 2015) and a batch size of $8$.